OpenStudy (anonymous):

Someone please help! ): F(x)= x+5 and G(x)= 11x-5 what is F(2)+G(-1) A. -9 B. -16 C. 14 D. 21

OpenStudy (anonymous):

F(2) = 2 + 5 = 7 G(-1) = 11 * -1 - 5 = -16 7 + -.16 = -9

OpenStudy (anonymous):

omg thank you would you be able to help me with more if i am stuck @moneybird